# DeepGB-TB: A Risk-Balanced Cross-Attention Gradient-Boosted Convolutional Network for Rapid, Interpretable Tuberculosis Screening

### v1

@@ Abstract
footnotetext: Corresponding authors.
1 Introduction
Tuberculosis (TB) inflicts a staggering toll on global health, remaining a leading infectious cause of death worldwide (World Health Organization 2023, 2024). The cornerstone of TB control, early diagnosis and treatment, is severely undermined in many regions by the limitations of conventional methods. Diagnostic tools like sputum smear microscopy and nucleic acid amplification tests (NAATs) either suffer from low sensitivity or are prohibitively expensive and require centralized laboratories and skilled technicians, creating significant barriers to access in low-resource settings (Lawn and Zumla 2011; Boehme et al. 2010). This diagnostic gap leads to delayed treatment, increased transmission, and preventable mortality (World Health Organization 2015), underscoring an urgent need for accessible, affordable, and scalable screening solutions.
The ubiquity of mobile phones presents a unique opportunity to bridge this gap. Cough, a cardinal symptom of pulmonary TB, contains a wealth of acoustic information that, if properly analyzed, could serve as a non-invasive digital biomarker (Imran, Posokhow et al. 2019). While prior research has explored AI for cough-based analysis, existing models often face critical limitations. Many either rely solely on audio, ignoring crucial demographic and clinical risk factors (e.g., age, sex, exposure history), or they struggle to effectively integrate these heterogeneous data types. Simple concatenation or late-stage fusion of features often fails to capture the complex, non-linear interplay between a patient’s background risk and their real-time acoustic symptoms, a process central to a clinician’s diagnostic reasoning (Ramachandram and Taylor 2017).
@@ 1 Introduction
2 Related Works
Early work in AI-driven TB diagnosis predominantly focused on analyzing Chest X-rays (CXRs), where Convolutional Neural Networks (CNNs) have achieved radiologist-level performance (Lakhani and Sundaram 2017). However, the reliance on specialized hardware for CXR imaging limits its applicability for widespread, community-level screening. This has motivated a research shift towards more accessible biomarkers, with cough sounds emerging as a leading alternative. The analysis of cough sounds via AI has emerged as a promising, low-cost screening paradigm. Initial approaches relied on traditional audio features paired with classical machine learning models like Logistic Regression (Cox 1958). More recently, deep learning has become the standard, with architectures such as 1D-CNNs (Kiranyaz, Ince, and Gabbouj 2016) and deeper models like ResNet (He et al. 2016) applied to spectrograms. The state-of-the-art (SOTA) is represented by pre-trained audio foundation models, notably Google’s Health Acoustic Representations (HeAR) (Google Research 2024), which has demonstrated superior performance across 33 health-related acoustic tasks, including TB recognition. Nevertheless, HeAR is not open-source and is only accessible via a restricted API for online inference. This creates a significant barrier to its adoption in resource-constrained settings and leaves a critical gap for an efficient, open, and specialized tool for TB screening. The clinical diagnosis of TB is inherently multimodal, integrating symptoms, patient history, and risk factors. In AI research, a common baseline for multimodal fusion is a late-fusion ensemble, such as a CNN-LightGBM model (Lu 2023), which often fails to capture deep inter-modal synergies. While advanced models for tabular data like the TabTransformer (Huang et al. 2020) exist, their effective integration with other modalities remains a challenge. At the frontier are Large Multimodal Models (LMMs) like Qwen-Omni (Xu et al. 2025), which represent general-purpose understanding. However, these models are computationally prohibitive and their superiority over smaller, specialized models for a specific diagnostic task is not yet established. Furthermore, the poor interpretability of many deep learning models (Zhang et al. 2021; Smith, Doe et al. 2022) remains a obstacle to their reliable clinical deployment.
3 Methodology
@@
3.1 Dataset
The dataset used in this study (Jaganath et al. 2024) was derived from a retrospective, multicenter case-control investigation encompassing 1,105 adult participants across seven countries (India, the Philippines, South Africa, Uganda, Vietnam, Tanzania, and Madagascar). Each participant presented with a new or worsening cough lasting at least two weeks, and for every subject, comprehensive clinical and demographic data were collected alongside cough acoustic recordings. Participants were classified as TB positive if they tested positive on any of the following microbiological reference assays: the Mycobacterial culture (Kent and Kubica 1985), the Cepheid Xpert MTB/RIF assay (Boehme et al. 2010), or Cepheid Xpert MTB/RIF Ultra assay (Chakravorty et al. 2017). TB negative participants tested negative on all assays. Ethical approval for this multicenter study was obtained from the institutional review boards of the participating centers. Exploratory analysis revealed significant group differences: TB prevalence was higher in males (33.2%) than females (19.7%); hemoptysis (43.2% vs. 24.3%), night sweats (39.0% vs. 17.3%), fever (40.0% vs. 16.1%), and weight loss (36.5% vs. 11.9%) were strongly associated with TB, motivating a multimodal approach.
3.2 Data Pre-processing and Feature Extraction
@@ 3.3 Statistical Analysis
A logistic regression model is then trained on the resulting embeddings to assess predictive power. We validate the structural integrity of this non-linear mapping using a Mantel correlation test (Mantel 1967; Wattenberg, Viégas, and Johnson 2016) and confirm the significance of the most discriminative features using Wald and independent two-sample t-tests.
3.4 Proposed Framework
@@
$f_{a,i}=\text{Flatten}\bigl{(}F^{(3)}_{i}\bigr{)},\quad$ (8)
@@
where $\hat{y}_{i}\in\mathbb{R}^{2}$ represents the predicted probabilities for TB-positive. The IMDM thus integrates tabular and audio features in a unified architecture, leveraging the strengths of boosting-based models for tabular data and CNNs for audio analysis, culminating in a robust, multimodal diagnostic framework.
For a fair comparison, all models undergo identical training, validation, and test splits. The training loss function, cross entropy loss (Goodfellow, Bengio, and Courville 2016), was maintained consistently throughout the experiments in deep learning models and was evaluated using metrics such as accuracy, sensitivity (TPR), specificity (TNR), and Area Under the Receiver Operating Characteristic Curve (AUROC) (Fawcett 2006). Statistical tests and a stepwise backward elimination approach (Babyak 2001) were employed to assess feature importance.
@@ 4 Experiments
4.1 Experimental Setup
We employed a 5-fold stratified cross-validation scheme for robust evaluation, which is ideal for the imbalanced nature of the dataset. All models were implemented in PyTorch 2.6.0 with CUDA 12.4 and Python 3.10, accelerated by an NVIDIA A100 80G GPU. We used the AdamW optimizer (Loshchilov and Hutter 2019) with a batch size of 64 and a cosine annealing learning rate schedule (minimum value of $1\times 10^{-6}$). Training durations were set by model type to ensure a fair comparison. Our proposed DeepGB-TB, along with other models trained from scratch (1D-CNN, ResNet, TabTransformer, and the CNN-LightGBM Ensemble), were trained for 100 epochs. Following standard fine-tuning practices, the large pre-trained models were trained for significantly fewer epochs: Wav2vec-2.0 Large for 3 epochs and the Qwen-Omni 3B with LoRA (Hu et al. 2021) for a single epoch.
@@ 4.1 Experimental Setup
✗ (Baseline) 0.765±0.0650.765$\pm$ 0.0650.765 ± 0.065 0.944±0.0530.944$\pm$ 0.0530.944 ± 0.053 0.567±0.1290.567$\pm$ 0.1290.567 ± 0.129 0.837±0.0800.837$\pm$ 0.0800.837 ± 0.080 0.887±0.0340.887$\pm$ 0.0340.887 ± 0.034
Gender 0.759±0.0440.759$\pm$ 0.0440.759 ± 0.044 0.806±0.0680.806$\pm$ 0.0680.806 ± 0.068 0.786±0.1310.786$\pm$ 0.1310.786 ± 0.131 0.796±0.0930.796$\pm$ 0.0930.796 ± 0.093 0.888±0.0240.888$\pm$ 0.0240.888 ± 0.024 0.82
Hemoptysis 0.764±0.0410.764$\pm$ 0.0410.764 ± 0.041 0.782±0.1160.782$\pm$ 0.1160.782 ± 0.116 0.761±0.1610.761$\pm$ 0.1610.761 ± 0.161 0.777±0.1250.777$\pm$ 0.1250.777 ± 0.125 0.885±0.0340.885$\pm$ 0.0340.885 ± 0.034 0.41
Weight Loss 0.770±0.0270.770$\pm$ 0.0270.770 ± 0.027 0.739±0.0780.739$\pm$ 0.0780.739 ± 0.078 0.622±0.1930.622$\pm$ 0.1930.622 ± 0.193 0.722±0.1300.722$\pm$ 0.1300.722 ± 0.130 0.881±0.0270.881$\pm$ 0.0270.881 ± 0.027 0.05
Smoke 0.779±0.0280.779$\pm$ 0.0280.779 ± 0.028 0.819±0.0590.819$\pm$ 0.0590.819 ± 0.059 0.766±0.0810.766$\pm$ 0.0810.766 ± 0.081 0.805±0.0720.805$\pm$ 0.0720.805 ± 0.072 0.888±0.0310.888$\pm$ 0.0310.888 ± 0.031 0.87
Fever 0.755±0.0390.755$\pm$ 0.0390.755 ± 0.039 0.856±0.0520.856$\pm$ 0.0520.856 ± 0.052 0.677±0.0400.677$\pm$ 0.0400.677 ± 0.040 0.807±0.0440.807$\pm$ 0.0440.807 ± 0.044 0.888±0.0350.888$\pm$ 0.0350.888 ± 0.035 0.91
Night Sweats 0.779±0.0200.779$\pm$ 0.0200.779 ± 0.020 0.806±0.0530.806$\pm$ 0.0530.806 ± 0.053 0.792±0.1250.792$\pm$ 0.1250.792 ± 0.125 0.798±0.0830.798$\pm$ 0.0830.798 ± 0.083 0.894±0.0320.894$\pm$ 0.0320.894 ± 0.032 0.21
Age 0.753±0.0290.753$\pm$ 0.0290.753 ± 0.029 0.836±0.0830.836$\pm$ 0.0830.836 ± 0.083 0.729±0.0780.729$\pm$ 0.0780.729 ± 0.078 0.807±0.0720.807$\pm$ 0.0720.807 ± 0.072 0.891±0.0310.891$\pm$ 0.0310.891 ± 0.031 0.57
Height 0.776±0.0140.776$\pm$ 0.0140.776 ± 0.014 0.748±0.1580.748$\pm$ 0.1580.748 ± 0.158 0.741±0.0840.741$\pm$ 0.0840.741 ± 0.084 0.754±0.1240.754$\pm$ 0.1240.754 ± 0.124 0.888±0.0310.888$\pm$ 0.0310.888 ± 0.031 0.87
Weight 0.759±0.0300.759$\pm$ 0.0300.759 ± 0.030 0.843±0.0860.843$\pm$ 0.0860.843 ± 0.086 0.525±0.2660.525$\pm$ 0.2660.525 ± 0.266 0.780±0.1570.780$\pm$ 0.1570.780 ± 0.157 0.886±0.0350.886$\pm$ 0.0350.886 ± 0.035 0.56
Cough Duration 0.754±0.0460.754$\pm$ 0.0460.754 ± 0.046 0.833±0.0560.833$\pm$ 0.0560.833 ± 0.056 0.735±0.0650.735$\pm$ 0.0650.735 ± 0.065 0.807±0.0630.807$\pm$ 0.0630.807 ± 0.063 0.883±0.0340.883$\pm$ 0.0340.883 ± 0.034 0.21
Heart Rate 0.764±0.0130.764$\pm$ 0.0130.764 ± 0.013 0.776±0.1570.776$\pm$ 0.1570.776 ± 0.157 0.724±0.1970.724$\pm$ 0.1970.724 ± 0.197 0.767±0.1540.767$\pm$ 0.1540.767 ± 0.154 0.882±0.0290.882$\pm$ 0.0290.882 ± 0.029 0.01
Temperature 0.764±0.0310.764$\pm$ 0.0310.764 ± 0.031 0.802±0.1030.802$\pm$ 0.1030.802 ± 0.103 0.666±0.0920.666$\pm$ 0.0920.666 ± 0.092 0.776±0.0930.776$\pm$ 0.0930.776 ± 0.093 0.885±0.0380.885$\pm$ 0.0380.885 ± 0.038 0.45
ZCR 0.790±0.0290.790$\pm$ 0.0290.790 ± 0.029 0.773±0.0590.773$\pm$ 0.0590.773 ± 0.059 0.804±0.1600.804$\pm$ 0.1600.804 ± 0.160 0.783±0.1080.783$\pm$ 0.1080.783 ± 0.108 0.890±0.0290.890$\pm$ 0.0290.890 ± 0.029 0.74
Centroid 0.800±0.0290.800$\pm$ $0$ 0.029 0.756±0.1180.756$\pm$ 0.1180.756 ± 0.118 0.679±0.0480.679$\pm$ 0.0480.679 ± 0.048 0.745±0.0890.745$\pm$ 0.0890.745 ± 0.089 0.886±0.0310.886$\pm$ 0.0310.886 ± 0.031 0.52
F0 0.762±0.0180.762$\pm$ 0.0180.762 ± 0.018 0.793±0.0610.793$\pm$ 0.0610.793 ± 0.061 0.607±0.1880.607$\pm$ 0.1880.607 ± 0.188 0.751±0.1160.751$\pm$ 0.1160.751 ± 0.116 0.889±0.0340.889$\pm$ 0.0340.889 ± 0.034 0.92
Energy 0.798±0.0300.798$\pm$ 0.0300.798 ± 0.030 0.740±0.1140.740$\pm$ 0.1140.740 ± 0.114 0.552±0.2790.552$\pm$ 0.2790.552 ± 0.279 0.707±0.1600.707$\pm$ 0.1600.707 ± 0.160 0.885±0.0360.885$\pm$ 0.0360.885 ± 0.036 0.43
Chroma Vector 0.792±0.0250.792$\pm$ 0.0250.792 ± 0.025 0.705±0.0880.705$\pm$ 0.0880.705 ± 0.088 0.581±0.2800.581$\pm$ 0.2800.581 ± 0.280 0.687±0.1650.687$\pm$ 0.1650.687 ± 0.165 0.890±0.0290.890$\pm$ 0.0290.890 ± 0.029 0.74
MFCCs 0.787±0.0270.787$\pm$ 0.0270.787 ± 0.027 0.803±0.0740.803$\pm$ 0.0740.803 ± 0.074 0.671±0.3480.671$\pm$ 0.3480.671 ± 0.348 0.777±0.1820.777$\pm$ 0.1820.777 ± 0.182 0.879±0.0410.879$\pm$ 0.0410.879 ± 0.041 0.05
Mel-Spectrogram 0.802±0.0350.802$\pm$ 0.0350.802 ± 0.035 0.763±0.0460.763$\pm$ 0.0460.763 ± 0.046 0.870±0.1160.870$\pm$ 0.1160.870 ± 0.116 0.790±0.0990.790$\pm$ 0.0990.790 ± 0.099 0.902±0.0240.902$\pm$ 0.0240.902 ± 0.024 $<$0.01
Table 2: Model performance comparison of DeepGB-TB with excluded features. * indicate independent samples t-test for AUROC compared to the full-variable (baseline); Data are mean $\pm$±SD. Bold values indicate statistical significance ($\leq$0.05).
4.2 Comparative Analysis
@@
We evaluated all models using a comprehensive set of metrics: Accuracy, True Positive Rate (TPR, or Sensitivity), True Negative Rate (TNR, or Specificity), F1-score, and Area Under the Receiver Operating Characteristic Curve (AUROC). As detailed in Table 1, our proposed DeepGB-TB achieves a SOTA AUROC of 0.903 and an F1-score of 0.851. This performance surpasses a wide range of baselines. For instance, it outperforms specialized unimodal models like TabTransformer (AUROC 0.737) and audio-only foundation models like HeAR (AUROC 0.768). It also exceeds the performance of a standard late-fusion CNN-LightGBM Ensemble (AUROC 0.792), highlighting the benefits of our integrated fusion architecture. Notably, while the large foundation model Qwen-Omni 3B demonstrates exceptional capability in the audio-only modality (AUROC 0.885), our more lightweight, specialized model still achieves superior overall performance when both modalities are integrated (0.903 compared to 0.900). This result further confirm that both modalities are crucial for our model’s success. The audio-only version of DeepGB-TB achieves an AUROC of 0.825, while the tabular-only version reaches 0.840, both significantly lower than the fully integrated model. This demonstrates our framework’s ability to seamlessly integrate cough acoustics and demographic data to enhance TB diagnosis, and its robust predictive performance reinforces its clinical reliability.
@@
4.3 Ablation Studies
To dissect the contribution of each component within the DeepGB-TB framework, we conduct a comprehensive ablation study to validate the contribution of each proposed module, with results detailed in Table 4. Starting with a simple CNN-Backbone, which yielded an AUROC of 0.809, we incrementally added our proposed modules. The integration of the CVPEM for tabular data improved the AUROC by 1.6% to 0.822. Subsequently adding the IMDM provided a substantial boost of 9.9% to 0.889. The further inclusion of our CM-BCA module advanced the AUROC to 0.901 (+11.4% over baseline). Finally, applying the TRBL brought the model to its peak performance of 0.903 AUROC and 0.851 F1-score. Furthermore, removing key components from the final DeepGB-TB model leads to a significant performance drop, with AUROC decreasing by 1.3% when CM-BCA is excluded. This step-wise improvement clearly demonstrates the synergistic effect and individual importance of each architectural innovation. We also investigated the optimal hyperparameter $\lambda$ for the TRBL function, which controls the penalty weight for false negatives. As shown in Table 4, we tested $\lambda$ values from 1 to 5. The model’s performance peaked at $\lambda=3$, achieving the highest AUROC (0.903) and F1-score (0.851). This finding confirms that a carefully calibrated, risk-sensitive loss function is critical for optimizing the model for the clinical priority of minimizing missed TB cases. On-device inference performance of our final quantized model (FP16) were evaluated using the ONNX Runtime. As presented in Table 5, the model demonstrates strong real-time capabilities across a range of edge devices. Notably, it achieves an inference latency of 142 ms on an iPhone 14 Pro while maintaining a high AUROC of 0.903. These results confirm the model’s efficiency and feasibility for practical, real-world deployment on resource-constrained hardware.
@@ 4.3 Ablation Studies
4.4 Feature Importance Analysis
To evaluate the contribution of individual characteristics to the predictive power of the model, we performed a leave-one-out feature analysis, the results summarized in Table 2. In this analysis, the model was trained and evaluated with one feature excluded at a time, and the resulting performance was compared to the baseline model using all features. A statistically significant drop in AUROC upon a feature’s exclusion indicates its importance. The results show that both clinical and acoustic features are vital for accurate diagnosis. The exclusion of Mel-Spectrogram features led to the most significant performance degradation (p$<$0.01), underscoring the critical role of comprehensive spectral representation. Among clinical symptoms, the absence of Hemoptysis (p=0.05), Weight Loss (p=0.05), and Heart Rate (p=0.01) each resulted in a significant decline in performance. Similarly, excluding MFCCs also significantly impacted the AUROC (p=0.05). The color intensity represents the magnitude of attention weights in Figure 4, indicating feature importance for the model’s prediction. For high-dimensional features, mean attention weights are computed across their respective dimensions. The x-axis corresponds to samples within the batch. The model consistently assigns higher weights to specific audio features (MFCCs, Mel-Spectrogram) and physiological signals (Heart Rate, Weight Loss), demonstrating its ability to identify and prioritize the most discriminative features from the multimodal input. This analysis validates our multimodal model effectively leverages a combination of high-level acoustic patterns and key clinical indicators to make its predictions.
@@ 4.4 Feature Importance Analysis
5 Conclusion and Limitations

### v2

@@ Abstract
Introduction
Tuberculosis (TB) inflicts a staggering toll on global health, remaining a leading infectious cause of death worldwide (Bagcchi 2023; Chen et al. 2025). The cornerstone of TB control, early diagnosis and treatment, is severely undermined in many regions by the limitations of conventional methods. Diagnostic tools like sputum smear microscopy and nucleic acid amplification tests (NAATs) either suffer from low sensitivity or are prohibitively expensive and require centralized laboratories and skilled technicians, creating significant barriers to access in low-resource settings (Snider Jr and Roper 1992; Boehme et al. 2010). This diagnostic gap leads to delayed treatment, increased transmission, and preventable mortality (Uplekar et al. 2015), underscoring an urgent need for accessible, affordable, and scalable screening solutions.
The ubiquity of mobile phones presents a unique opportunity to bridge this gap. Cough, a cardinal symptom of pulmonary TB, contains a wealth of acoustic information that, if properly analyzed, could serve as a non-invasive digital biomarker (Kant and Srivastava 2018). While prior research has explored AI for cough-based analysis, existing models often face critical limitations. Many either rely solely on audio, ignoring crucial demographic and clinical risk factors (e.g., age, sex, exposure history), or they struggle to effectively integrate these heterogeneous data types. Simple concatenation or late-stage fusion of features often fails to capture the complex, non-linear interplay between a patient’s background risk and their real-time acoustic symptoms, a process central to a clinician’s diagnostic reasoning (Jabeen et al. 2023).
@@ Introduction
Related Works
Early work in AI-driven TB diagnosis predominantly focused on analyzing Chest X-rays (CXRs), where Convolutional Neural Networks (CNNs) have achieved radiologist-level performance (Lakhani and Sundaram 2017). However, the reliance on specialized hardware for CXR imaging limits its applicability for widespread, community-level screening. This has motivated a research shift towards more accessible biomarkers, with cough sounds emerging as a leading alternative. The analysis of cough sounds via AI has emerged as a promising, low-cost screening paradigm. Initial approaches relied on traditional audio features paired with classical machine learning models like Logistic Regression (Cox 1958). More recently, deep learning has become the standard, with architectures such as 1D-CNNs (Kiranyaz, Ince, and Gabbouj 2016) and deeper models like ResNet (He et al. 2016) applied to spectrograms. The state-of-the-art (SOTA) includes pre-trained audio foundation models like Google’s HeAR (Baur et al. 2024), which demonstrated superior performance on 33 health-related acoustic tasks, including TB recognition. Nevertheless, HeAR is not open-source and accessible only via a restricted API for online inference, limiting adoption in resource-constrained settings and creating a critical gap for an efficient, open-source TB screening tool. Clinical TB diagnosis is inherently multimodal, integrating symptoms, history, and risk factors. Common multimodal fusion baselines, such as late-fusion CNN-LightGBM ensembles (Lu 2023), often fail to capture deep inter-modal synergies (Li et al. 2025a). While advanced tabular models like TabTransformer (Huang et al. 2020) exist, their effective integration remains challenging. Frontier models include Large Multimodal Models (LMMs) like Qwen-Omni (Xu et al. 2025), offering general-purpose understanding (Li et al. 2025b; Zhang et al. 2025). However, LMMs are computationally prohibitive, and their superiority over specialized models for this diagnostic task is unestablished. Furthermore, the poor interpretability of many deep learning models (Al-Zoghby et al. 2025; Sathishkumar et al. 2024) remains an obstacle to reliable clinical deployment.
Methodology
@@
Dataset
The dataset used in this study (Jaganath et al. 2024) was derived from a retrospective, multicenter case-control investigation encompassing 1,105 adult participants across seven countries (India, the Philippines, South Africa, Uganda, Vietnam, Tanzania, and Madagascar). Each participant presented with a new or worsening cough lasting at least two weeks. Comprehensive clinical, demographic, and cough acoustic data were collected for each participant. Participants were classified as TB positive if testing positive on Mycobacterial culture (Kent 1985), Xpert MTB/RIF (Boehme et al. 2010), or Xpert Ultra (Chakravorty et al. 2017), and TB negative if negative on all assays. The study received ethical approval from all participating institutional review boards. Exploratory analysis revealed significant group differences: TB prevalence was higher in males (33.2%) than females (19.7%). Furthermore, hemoptysis (43.2% vs. 24.3%), night sweats (39.0% vs. 17.3%), fever (40.0% vs. 16.1%), and weight loss (36.5% vs. 11.9%) were all strongly associated with TB.
Data Pre-processing and Feature Extraction
@@
A logistic regression model is then trained on the resulting embeddings to assess predictive power. We validate the structural integrity of this non-linear mapping using a Mantel correlation test (Mantel 1967; Linderman and Steinerberger 2019) and confirm the significance of the most discriminative features using Wald and independent two-sample t-tests.
Proposed Framework
@@
$f_{a,i}=\text{Flatten}\bigl(F^{(3)}_{i}\bigr)$$\quad$ (8)
@@
where $\hat{y}_{i}\in\mathbb{R}^{2}$ represents the predicted probabilities for TB-positive. The IMDM thus integrates tabular and audio features in a unified architecture, leveraging the strengths of boosting-based models for tabular data and CNNs for audio analysis, culminating in a robust, multimodal diagnostic framework. For a fair comparison, all models undergo identical training, validation, and test splits. The training loss function, cross entropy loss (Goodfellow et al. 2016), was maintained consistently throughout the experiments in deep learning models and was evaluated using metrics such as accuracy, sensitivity (TPR), specificity (TNR), and Area Under the Receiver Operating Characteristic Curve (AUROC) (Fawcett 2006). Statistical tests and a stepwise backward elimination approach (Babyak 2004) were employed to assess feature importance.
@@
Experimental Setup
We employed a 5-fold stratified cross-validation scheme for robust evaluation, which is ideal for the imbalanced nature of the dataset. All models were implemented in PyTorch 2.6.0 with CUDA 12.4 and Python 3.10, accelerated by an NVIDIA A100 80G GPU. We used the AdamW optimizer (Loshchilov and Hutter 2019) with a batch size of 64 and a cosine annealing learning rate schedule (minimum value of $1\times 10^{-6}$). Training durations were set by model type to ensure a fair comparison. Our proposed DeepGB-TB, along with other models trained from scratch (1D-CNN, ResNet, TabTransformer, and the CNN-LightGBM Ensemble), were trained for 100 epochs. Following standard fine-tuning protocols, the large pre-trained models were trained for limited epochs: Wav2vec-2.0 for 3 epochs and Qwen-Omni with LoRA (Hu et al. 2021) for a single epoch.
@@
✗ (Baseline) $0.765\pm 0.065$ $0.944\pm 0.053$ $0.567\pm 0.129$ $0.837\pm 0.080$ $0.887\pm 0.034$
Gender $0.759\pm 0.044$ $0.806\pm 0.068$ $0.786\pm 0.131$ $0.796\pm 0.093$ $0.888\pm 0.024$ 0.82
Hemoptysis $0.764\pm 0.041$ $0.782\pm 0.116$ $0.761\pm 0.161$ $0.777\pm 0.125$ $0.885\pm 0.034$ 0.41
Weight Loss $0.770\pm 0.027$ $0.739\pm 0.078$ $0.622\pm 0.193$ $0.722\pm 0.130$ $881\pm 0.027 0.0$5
Smoke $0.779\pm 0.028$ $0.819\pm 0.059$ $0.766\pm 0.081$ $0.805\pm 0.072$ $0.888\pm 0.031$ 0.87
Fever $0.755\pm 0.039$ $0.856\pm 0.052$ $0.677\pm 0.040$ $0.807\pm 0.044$ $0.888\pm 0.035$ 0.91
Night Sweats $0.779\pm 0.020$ $0.806\pm 0.053$ $0.792\pm 0.125$ $0.798\pm 0.083$ $0.894\pm 0.032$ 0.21
Age $0.753\pm 0.029$ $0.836\pm 0.083$ $0.729\pm 0.078$ $0.807\pm 0.072$ $0.891\pm 0.031$ 0.57
Height $0.776\pm 0.014$ $0.748\pm 0.158$ $0.741\pm 0.084$ $0.754\pm 0.124$ $0.888\pm 0.031$ 0.87
Weight $0.759\pm 0.030$ $0.843\pm 0.086$ $0.525\pm 0.266$ $0.780\pm 0.157$ $0.886\pm 0.035$ 0.56
Cough Duration $0.754\pm 0.046$ $0.833\pm 0.056$ $0.735\pm 0.065$ $0.807\pm 0.063$ $0.883\pm 0.034$ 0.21
Heart Rate $0.764\pm 0.013$ $0.776\pm 0.157$ $0.724\pm 0.197$ $0.767\pm 0.154$ $882\pm 0.029 0.0$1
Temperature $0.764\pm 0.031$ $0.802\pm 0.103$ $0.666\pm 0.092$ $0.776\pm 0.093$ $0.885\pm 0.038$ 0.45
ZCR $0.790\pm 0.029$ $0.773\pm 0.059$ $0.804\pm 0.160$ $0.783\pm 0.108$ $0.890\pm 0.029$ 0.74
Centroid $0.800\pm 0.029$ $0.756\pm 0.118$ $0.679\pm 0.048$ $0.745\pm 0.089$ $0.886\pm 0.031$ 0.52
F0 $0.762\pm 0.018$ $0.793\pm 0.061$ $0.607\pm 0.188$ $0.751\pm 0.116$ $0.889\pm 0.034$ 0.92
Energy $0.798\pm 0.030$ $0.740\pm 0.114$ $0.552\pm 0.279$ $0.707\pm 0.160$ $0.885\pm 0.036$ 0.43
Chroma Vector $0.792\pm 0.025$ $0.705\pm 0.088$ $0.581\pm 0.280$ $0.687\pm 0.165$ $0.890\pm 0.029$ 0.74
MFCCs $0.787\pm 0.027$ $0.803\pm 0.074$ $0.671\pm 0.348$ $0.777\pm 0.182$ $879\pm 0.041 0.0$5
Mel-Spectrogram $0.802\pm 0.035$ $0.763\pm 0.046$ $0.870\pm 0.116$ $0.790\pm 0.099$ $902\pm 0.024 <<0.0$1
[…]
Comparative Analysis
@@
We evaluated all models using a comprehensive set of metrics: Accuracy, True Positive Rate (TPR, or Sensitivity), True Negative Rate (TNR, or Specificity), F1-score, and Area Under the Receiver Operating Characteristic Curve (AUROC). As detailed in Table 1, our proposed DeepGB-TB achieves SOTA performance (0.903 AUROC, 0.851 F1-score). This surpasses unimodal baselines like TabTransformer (0.737 AUROC) and audio-only foundation models like HeAR (0.768 AUROC), as well as a late-fusion CNN-LightGBM Ensemble (0.792 AUROC), highlighting our integrated architecture’s superiority. Notably, our model achieves superior multimodal performance over the larger Qwen-Omni 3B (0.903 vs. 0.900 AUROC), despite the latter’s strong audio-only result (0.885 AUROC). The framework’s effective synergy is confirmed by internal ablations: the audio-only (0.825 AUROC) and tabular-only (0.840 AUROC) versions perform significantly worse than the fully integrated model, demonstrating our framework’s effective synergy between cough acoustics and demographic data to enhance TB diagnosis, and its robust predictive performance reinforces its clinical reliability.
@@
Ablation Studies
We conducted a comprehensive ablation study to quantify the contribution of each module, with results in Table 3. Starting from a CNN-Backbone baseline (0.809 AUROC), we incrementally added components: the CVPEM improved AUROC to 0.822 (+1.6%), subsequently adding the IMDM substantially boosted it to 0.889 (+9.9%), and the CM-BCA module further advanced it to 0.901 (+11.4%). Applying the TRBL achieved peak performance (0.903 AUROC). This step-wise improvement demonstrates component synergy; conversely, removing modules like CM-BCA from the final model incurs a significant 1.3% AUROC drop. We also optimized the TRBL hyperparameter $\lambda$, which weights the false negative penalty. As shown in Table 4, performance peaked at $\lambda=3$ (0.903 AUROC) when testing values from 1 to 5, confirming that a calibrated, risk-sensitive loss is critical for the clinical priority of minimizing missed TB cases. Finally, we evaluated the on-device inference of the FP16-quantized model using ONNX Runtime (Table 5). The model exhibits strong real-time capabilities, achieving a 142 ms latency on an iPhone 14 Pro while maintaining 0.903 AUROC, confirming its efficiency and feasibility for deployment on resource-constrained hardware.
@@
Feature Importance Analysis
To evaluate feature importance, we performed a leave-one-out analysis (Table 2), iteratively training the model with one feature excluded and comparing its AUROC to the full-feature baseline. A statistically significant performance drop (p<<0.05) indicates feature importance. Results show that both clinical and acoustic features are vital. Excluding Mel-Spectrograms caused the most significant degradation (p<<0.01), underscoring the critical role of spectral representation. Among clinical data, excluding Hemoptysis (p=0.05), Weight Loss (p=0.05), or Heart Rate (p=0.01) significantly reduced performance, as did excluding MFCCs (p=0.05). We further visualize feature importance via attention weights in Figure 4, where color intensity reflects magnitude (mean-computed for high-dimensional features) and the x-axis denotes batch samples. The model consistently assigns higher weights to key audio features (MFCCs, Mel-Spectrogram) and clinical signals (Heart Rate, Weight Loss), demonstrating its ability to prioritize discriminative inputs. Both analyses validate that our multimodal model effectively leverages a combination of high-level acoustic patterns and key clinical indicators for its predictions.
Conclusion and Limitations
@@
Acknowledgments
The datasets used for the analyses described were contributed by Dr. Adithya Cattamanchi at UCSF and Dr. Simon Grandjean Lapierre at University of Montreal and were generated in collaboration with researchers at Stellenbosch University (PI Grant Theron), Walimu (PIs William Worodria and Alfred Andama); De La Salle Medical and Health Sciences Institute (PI Charles Yu), Vietnam National Tuberculosis Program (PI Nguyen Viet Nhung), Christian Medical College (PI DJ Christopher), Centre Infectiologie Charles Mérieux Madagascar (PIs Mihaja Raberahona & Rivonirina Rakotoarivelo), and Ifakara Health Institute (PIs Issa Lyimo & Omar Lweno) with funding from the U.S. National Institutes of Health (U01 AI152087), The Patrick J. McGovern Foundation and Global Health Labs.